OpenStudy (anonymous):

Find the slope of the line passing through the points (4, 1) and (2, 0).

OpenStudy (wasiqss):

2

OpenStudy (anonymous):

thanks this was kinda driving me nuts

OpenStudy (hoblos):

well it is (0-1)/(2-4) = -1/-2 = 1/2

OpenStudy (anonymous):

(0-1)/(2-4) -1/-2 1/2

OpenStudy (anonymous):

Slope m = difference of y / difference of x

OpenStudy (anonymous):

http://cs.selu.edu/~rbyrd/math/slope/

OpenStudy (anonymous):

0-1/2-4=1/2